OpenStudy (anonymous):

Solve 4(a-3)+5=7(3a-1)

OpenStudy (anonymous):

I get 4a=21a... But I dont think thats right...

OpenStudy (anonymous):

4a - 12 +5 = 21a - 7 4a - 7=21a -7 4a = 21a now i dont know where to go from here either.

OpenStudy (anonymous):

a=0

OpenStudy (anonymous):

Clever. Thanks! :)

OpenStudy (anonymous):

but how?

OpenStudy (anonymous):

Because that's the only way to make them both equal to each other.

OpenStudy (anonymous):

0 = 17a = 0? thanks, man i feel dumb sometimes.